OpenStudy (anonymous):

Find the sum of the following infinite geometric series, if it exists. 1/2+(-1/4)+1/8+(-1/16)+

OpenStudy (solomonzelman):

what is the common ratio ?

OpenStudy (anonymous):

-1/2

OpenStudy (solomonzelman):

yes`

OpenStudy (solomonzelman):

This is an infinite series correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is the answer 1/3

OpenStudy (solomonzelman):

Ok, so normally, as you know the sum of a geometric series with common ratio r, and first term \(a_1\) is given by: \(\large\color{black}{ \displaystyle {\rm S}=\frac{{\rm a}_1(1-{\rm r}^n)}{1-{\rm r}} }\) Now, for any infinite geometric series (where |r|<1, which is true in this case) \({\rm r}^n\) will be approaching 0, and thus \(1-{\rm r}^n\) will be equivalent to 1-0=1. \({\rm a}_1~\cdot 1\), is same as just \({\rm a}_1\). So for an infinite geometric series we therefore write: \(\large\color{black}{ \displaystyle {\rm S}=\frac{{\rm a}_1}{1-{\rm r}} }\)

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle {\rm S}=\frac{\frac{1}{2}}{1-\left(-\frac{1}{2}\right)} }\) \(\large\color{black}{ \displaystyle {\rm S}=\frac{\frac{1}{2}}{\frac{2}{2}-\left(-\frac{1}{2}\right)} }\) \(\large\color{black}{ \displaystyle {\rm S}=\frac{\frac{1}{2}}{\frac{2}{2}+\frac{1}{2}} }\) \(\large\color{black}{ \displaystyle {\rm S}=\frac{\frac{1}{2}}{\frac{3}{2}} }\) \(\large\color{black}{ \displaystyle {\rm S}=\frac{1}{3} }\) your answer is correct!!

OpenStudy (anonymous):

yay THANKS!!!!!

OpenStudy (solomonzelman):

yw